### 10 Mar 2026 v1

# Strong Electron Correlation Identified in Planetary Atomic Structure

Abstract
Investigating the two-electron correlation of doubly excited states (DESs) in the planetary atomic structure poses multiple conceptual and methodological challenges. Such highly correlated dynamics in a three-body Coulomb system can leave distinct signatures in the unique nonsequential above-threshold double ionization (NS-ATDI) exposed in a coherent laser field. Herein, we perform kinematically complete studies for multi-photon double ionization of cold strontium atoms. Our results reveal strongly favored NS-ATDI characterized by a series of band structures that demonstrate the energy and angular correlations of emitted two-electron pairs. These unprecedented characteristics reinforce the fact that both electrons are excited and ionized via the transition DESs, such that exotically strong two-electron correlations associated with atomic structure are relayed in the laser-driven time-dependent three-body system. The present findings transcend the traditional landscape in multi-photon double ionization and reshape our fundamental understanding of electron correlation in nature.
Electron correlation in Coulombic systems forms the cornerstone of many-body physics and governs the electronic energies and structures of complex microscopic and macroscopic systems. It is especially critical for strongly correlated materials including high-temperature superconductors (?) and colossal magneto resistance in transition-metal oxides (?). Studies on this topic often center on a prototypical platform provided by a helium or helium-like atom, where key phenomena including Fano resonance of doubly excited states (DESs) occur (?, ?, ?, ?, ?, ?, ?, ?). These are the simplest three-body systems in nature, but complex enough to challenge the most advanced calculations (?). However, understanding the dynamics behind correlated electrons has been far from sufficient up to now.
DESs exhibit strong electron correlations characterized by interdependent motion of two electrons in position and momentum space (?, ?). When two electrons in high-lying DESs orbit the nucleus in distinct paths, analogous to planets orbiting a star, the system constitutes a so-called planetary atom (?) —a strongly correlated three-body system that has attracted increasing interest for decades (?, ?, ?, ?, ?, ?). Ion-yield and absorption spectroscopy have served as standard experimental tools to investigate planetary atoms (?, ?, ?), but such studies have been largely limited to measuring the energies of DESs. Direct access to the underlying structural dynamics has remained elusive, even though classical Kepler planetary motions in the eZe and Zee configurations have been predicted to exhibit angular stability (?, ?, ?). In these simplified one-dimensional pictures, the eZe configuration corresponds to two electrons on opposite sides of the nucleus, whereas both electrons in the Zee configuration are on the same side. Remarkably, the angular correlation of the electrons is preserved from high-lying DESs into the continuum above the double ionization (DI) threshold, even across the quantum chaotic regime (?, ?, ?, ?).
DI serves as a sensitive probe of such correlated behavior, especially nonsequential DI, because the energy and momentum spectra of the two outgoing correlated electrons carry rich information that reveals the underlying physical mechanism. As an outstanding benchmark, single-photon DI studies have uncovered the canonical knock-out and shake-off ionization mechanisms. Investigations have been extended to two-photon (?, ?, ?, ?, ?), multi-photon (?, ?), and strong-field induced (?, ?, ?) DIs, revealing rich intrinsic ionization dynamics that remain active research frontiers. In the multi-photon strong perturbation regime, the above-threshold ionization phenomenon for two-electron emission—known as nonsequential above-threshold double ionization (NS-ATDI)—gives rise to highly correlated two-electron wave packets, in which both electrons absorb photons and share the excess energy in units of one photon. Using numerical solutions to the full-dimensional two-electron time-dependent Schrödinger equation (TDSE), Parker et al. (?) calculated ATDI for helium in high-coherence, high-brightness XUV laser fields and found a hint of nonsequential signals in the coexistence of primary sequential DI processes. There have been experimental reports on ATDI of rare gas atoms (?, ?) and molecules (?) induced by 400 nm lasers. To date, all experimental and theoretical studies indicate that sequential processes overwhelmingly dominate in ATDI.
Once high-lying DESs participate in NS-ATDI, a complex time-dependent three-body problem emerges, as the two-electron dynamics within the planetary structure evolves in the time-dependent laser field. The strong radial and angular correlations encoded in transient DESs (?, ?) can be accessed by detecting the emitted electron pairs, yielding distinct and revealing signatures in NS-ATDI. However, NS-ATDI studies involving DESs as intermediate transition states remain poorly developed because of substantial methodological challenges. Theoretically, single-photon perturbation theory and simple mean-field approaches are insufficient to describe this physical process. Furthermore, owing to multi-photon absorption, the coherence of the laser field is imprinted onto the highly correlated transition and emission processes, posing a great challenge for achieving a clear physical picture.
In this work, we use strontium atom—a helium-like system—as an ideal platform to investigate two-electron correlations. In this system, a large number of DESs can be efficiently populated via multi-photon excitation with optical lasers (?, ?, ?, ?, ?), owing to much larger excitation cross sections than those of helium with one-photon absorption. Specifically, we perform the first kinematically complete study on ATDI by exploiting a state-of-the-art magneto-optical trap reaction microscope (MOTReMi) (?, ?, ?) and numerically solving the full-dimensional two-electron TDSE (?, ?, ?). A schematic diagram of Sr-MOTReMi is shown in Fig. 1. The cold strontium atoms created by laser cooling technologies were doubly ionized by femtosecond lasers. The momenta of Sr2+ ion and two photoelectrons in three dimensions were detected in coincidence with extremely high resolutions. The joint energy spectra (JES) from DI of the ground state show significantly dominant NS-ATDI features where two electrons present band structures and strong angular correlations as equally sharing excess energy. As intermediate transition states, DESs relay the electron correlations (see in the inset of Fig. 1), thereby enabling these distinct NS-ATDI features in the laser-driven time-dependent three-body system.
ATDIs at 800 nm
In Fig. 2, A to C, we present the JES for DI of driven by a linearly polarized 800 nm laser at various laser intensities, and electron-electron joint momentum spectra (JMS) can be seen in the Supplementary Fig. S1. These results clearly demonstrate the multi-photon absorption in a strong perturbative regime. The total energy for each ATDI structure is given by , where is a photon energy, and represent the kinetic energies of the two photoelectrons , and = 5.7 eV/ = 11.0 eV denote the first/second ionization potential. The number of absorbed photons is estimated to be -. 0.4 eV is the ponderomotive shift at a laser peak intensity = 6 TW/cm2. For the sequential DI of , four and eight-photon are required for the ionization processes Sr Sr+ and Sr Sr2+, respectively. For the nonsequential DI of , at least 11 photons are required. However, the 11-photon channel is effectively closed due to the shift induced by .
Distinct NS-ATDI features are observed in the JES. Taking the 13-photon channel in Fig. 2B as an example, NS-ATDI shows a pronounced band pattern in the middle of the diagonal stripe, which is also characterized by circular arcs in the JMS (see Fig. S1 in the Supplementary). Two emitted electrons share a total energy of eV, i.e., the kinetic energy of one electron increases as the other one decreases or vice versa, keeping their total energy constant. In contrast, sequential ATDI appears as island structures, one of which is indicated in the green dashed circle in Fig. 2B. In this case, the two sequentially ionized electrons have different kinetic energies, with eV from the single ionization (SI) of Sr ( see also the peak I in Fig. 2D) and eV from the first order above-threshold ionization (ATI) of Sr+. As the laser intensity increases, these features of NS-ATDI always prevail, although the stripes start to become broad because of the varying over the entire laser pulse. The continuous distribution of these band structures constitutes a definitive hallmark of nonsequential ATDI and reflects two-electron energy correlations induced by multi-photon absorption. Most notably, the nonsequential signals are remarkably prominent, even comparable to those of sequential processes. These findings are quite different from previous theoretical and experimental ATDI studies (?, ?, ?), where sequential ATDIs always dominate absolutely. For sequential ATDI, two electrons are sequentially ionized from Sr and Sr+ and their kinetic energies depend on the number of absorbed photons and the ionization potentials Ip, resulting in the absence of energy correlations.
Strong electron correlations are further manifested in the correlated angular distribution (CAD) of the two photoelectrons. For nonsequential electron pairs with nearly equal kinetic energies eV located in the yellow dashed box in Fig. 2B, the CAD of one electron is plotted in Fig. 2E with the emission direction of the other fixed. The maximum ionization probability of the second electron appears at a relative emission angle , indicated by a red arrow. The underlying mechanisms for the almost back-to-back emission might refer to the planetary atomic structure of the transition DESs, where the classical eZe configuration reflecting angular correlation survives significantly at relatively large angles between two electrons for the principal series (?, ?). For the principal series of DESs, the angle can approach 180∘ as closely as possible when approaching the DI threshold (?, ?). The enormous density of DESs might lead to nontrivial mixing, so even qualitative analysis of the DESs can be difficult. Nevertheless, a detailed and complete understanding still challenges current many-body theories and needs further investigation. The CAD for two sequential electrons is also shown in Fig. 2F. As expected, we observe almost isotropic distributions, i.e., one electron is emitted independently on the other, and no maximum in CAD is present. Note that progressively decreasing counts at relatively small angles mainly results from the decreasing detection efficiency when two electrons arrive at the detector with approaching time. It has been confirmed that the CAD profile for NS-ATDI in Fig. 2E does not change after calibration with the data of sequential ATDI in Fig. 2F. In addition, the data in the regions very close to 0∘ and 180∘ are not physically meaningful due to the solid angles.
Considering the two-electron energy levels of strontium, at 800 nm, multi-photon resonant excitation to DESs should have been involved in the intermediate stage. Subsequently, the outermost two electrons in highly excited orbitals can be simultaneously ionized with nearly equal energy sharing and back-to-back emission. To verify the role of DESs, we plot the energy spectra of the SI in Fig. 2D at various laser intensities, where electrons are collected in coincidence with Sr+ in order to avoid contributions from DI. Obviously, the distribution of the peak III, i.e., the first ATI, gradually varies from being symmetric to being asymmetric as the laser intensity increases, indicating the onset of autoionization decay from DESs. Most importantly, the autoionization features within the dashed box coincide closely with the emergence of NS-ATDI: At TW/cm2, we find no DI signals, while the autoionization of DESs is also negligible in Fig. 2D. As autoionization signals gradually present at TW/cm2 and 18 TW/cm2, NS-ATDI structures accordingly become correspondingly prominent. However, it remains difficult to identify which specific DESs give rise to the spectral change near the first ATI because of the presence of a dense manifold of DESs. Nevertheless, for the first time, the above experimental observations present the scenario that dominating NS-ATDI features result from simultaneous emission of two electrons by coherent absorption of a couple of photons via one or a few DESs. The pronounced energy and angular correlations of two-electron pairs directly reflect the characteristic ionization dynamics of DESs.
Furthermore, we point out that the pathways to intermediate DESs can be manipulated by changing initial states and the ellipticity of the laser pulse, and subsequently modify the final JES results significantly. This assertion is supported by the ATDIs of the excited state of ), as shown in Fig. S2A of the Supplementary Materials. Interestingly, the JES presents visible and discrete island structures from sequential ATDIs, distinctly different from those in Fig. 2A. Changing the initial state from to will modify the character of intermediate transition states, which strongly perturbs electron correlations and ultimately suppresses the NS-ATDI channel. Meanwhile, by using a circularly polarized laser to ionize the atoms, sequential ATDIs with clear ‘islands’ structures cover the JES, as shown in Fig. S2B. This may be explained as follows: once a circularly polarized photon is absorbed, the quantum number of the total angular momentum prefers to increase by one, so that it is highly detrimental to the formation of DESs during the absorption of more than 12 photons. Consequently, the strong correlation is intercepted during the transition stages, and the electrons tend to be released one-by-one.
ATDIs at 400 nm
To the best of our knowledge, so far, there is no report on a rigorous calculation of multi-photon induced ATDI at 800 nm or longer wavelengths for heavy atoms with two outermost electrons. The extremely large computation demand prohibits the attempt under current theoretical methodologies.
To further consolidate the above finding about the strong NS-ATDI, we carried out additional and independent measurements for a linearly polarized laser at 400 nm, whose results can be directly compared with our ab initio calculations.
In this shorter wavelength region, the field-induced electron collision is further suppressed and potential pathways of resonant excitation are also simplified. The experimental JES together with the results of the two-electron TDSE calculation are shown in Fig. 3, where the zero and first orders of ATDI are observed at = 18 TW/cm2. We note that the nonsequential DI band in the experimental measurements remains the dominant structure located at eV, whose overall features are reproduced by the theoretical calculation. The simultaneous excitation of DESs during the multi-photon absorption, as we discussed before, is very likely a universal correlation mechanism that induces the strong signal in the NS-ATDI channel.
Conclusions
With a state-of-the-art MOTReMi and the numerical solution to the full-dimensional two electron TDSE, we performed kinematically complete studies for multi-photon-induced ATDI in a helium-like heavy atom—strontium. Our experimental results exhibit prominently dominant NS-ATDI features. The emitted electron pairs exhibit strong energy and angular correlations, directly revealing ionization dynamics of the transient DESs. SI electron spectra reveal clear contributions from DES autoionization, consistent with NS-ATDI. This observation provides direct evidence that DESs are involved in mediating strong electron correlation. These remarkable, previously unobserved nonsequential features arise from pathways in which tightly bound two electrons, which simultaneously absorb coherent photons, are excited and ionized via intermediate DES ladders. Our work represents a significant advance in understanding strong two-electron correlation in laser-driven time-dependent three-body Coulomb systems and offers new insights into electron correlation dynamics across microscopic and macroscopic quantum phenomena.
Materials and methods
Experimental methods
A schematic diagram of the experimental setup, the Sr-MOTReMi (?, ?), is shown in Fig. 1. The hot strontium atoms evaporated in an oven are decelerated by a Zeeman slower laser beam and then precooled and trapped by two orthogonal pairs of retroreflected two-dimensional (2D) cooling laser beams. Subsequently, the 2D MOT is propelled to generate a cold atomic beam by a pushing laser beam and further cooled and trapped by a three-dimensional (3D) MOT configuration. Consequently, cold atomic targets for our setup can be prepared in the 2D MOT beam, molasses beam, and 3D MOT. The 2D MOT beam delivers the continuous cold atomic beam employed for photoionization measurements of Sr(). When the six 3D cooling laser beams are turned on in the reaction chamber without the gradient magnetic field, the molasses target containing cold atoms in the Sr() state is prepared for measurements. The 3D MOT is used for the overlapping determination of laser focus and atomic beams. We used a Ti:sapphire laser pulses with 800 nm and 400 nm wavelengths and 50 fs pulse duration (FWHM) at intensities of 3-60 TW/cm2 to produce ions and electrons, which are collected in a 4 angle in a standard ReMi approach, see detailed configurations in Section 1 of Supplementary Materials. The momentum resolution for coincident Sr+/Sr2+ and electrons along the longitudinal direction of the spectrometer is achieved up to 0.12 a.u. and 0.02 a.u., respectively, and slightly lower in the transverse plane.
Theoretical methods
The DI of the Sr atom is simulated using state-of-the-art numerical methods for solving the TDSE (?, ?, ?). Here, we assume that only two electrons are active and interact directly with the external laser field. The effects of inner-shell electrons and the nuclear core are incorporated through an angular-momentum-dependent pseudopotential, for details see Section 4 in Supplementary Materials. The two-electron wave function is expanded in coupled spherical harmonics, while its radial components are discretized using the finite-element discrete variable representation (FE-DVR) method. Both the maximum total angular momentum quantum number and the maximum single-electron angular momentum quantum number are set to 7. Nonuniform finite elements are employed, with finer element sizes near the core region and progressively larger elements at greater distances. Time propagation of the two-electron wave function is performed using the Lanczos method. During propagation, the two electrons are confined within a spherical box of radius 250 a.u.. The wave function is periodically split into inner and outer components. The inner part is always propagated numerically under the full Hamiltonian, whereas the outer part is propagated analytically using the Volkov propagator. At the end of the propagation, DI signals are extracted from both the inner and outer wave-function components.
Acknowledgments
Funding:
This work is supported by the National Key Research and Development Program of China (Nos. 2022YFA1604302 and 2024YFA1612101), the National Natural Science Foundation of China (Nos. 12334011, 12234002, 12541505, 12204498, 12174284 and 12474346), and the Natural Science Foundation of Henan (Grant No. 252300421304). We also acknowledge support from the Shanghai-XFEL beamline project (SBP) (Grant No. 31011505505885920161A2101001) and the computational resource of the HPC platform in ShanghaiTech University.
Author contributions:
Y.J. supervised the project. X.Y., Z.S., S.R., J.L., Z.W., X.W., and Y.J. performed the experiments. Y.H., Y.F., W.J., and L.Y.P. performed the simulations. X.Y., Z.S., and Y.J. analyzed the data. Y.J., Z.S., and X.Y. wrote the manuscript after in-depth discussions with K.U., A.C., and with important input from all other authors.
Competing interests:
The authors have no competing interests to declare
Data and materials availability:
The data and materials that support the plots within this article are available from the corresponding authors upon reasonable request.